# Confidence-Guided Human-AI Collaboration: Reinforcement Learning with Distributional Proxy Value Propagation for Autonomous Driving

### v1

@@ II-D Overall Training Process
$J^{TD}_{\mathcal{Z}^{c}}(\zeta)=-\underset{\begin{subarray}{c}\left(s,a_{g},r,s^{\prime}\right)\sim\mathcal{B}_{g},a^{\prime}\sim\pi^{g}_{\bar{\phi}},\\ Z\left(s^{\prime},a^{\prime}\right)\sim\mathcal{Z}^{c}_{\bar{\zeta}}\left(\cdot\mid s^{\prime},a^{\prime}\right)\end{subarray}}{\mathbb{E}}\left[\log\mathcal{P}\left(y_{z}^{g}\mid\mathcal{Z}^{c}_{\zeta}(\cdot\mid s,a)\right)\right],$ (38)
@@
$y_{z}^{g}=r+\gamma\left(Z\left(s^{\prime},a^{\prime}\right)-\alpha\log\pi^{g}_{\bar{\phi}}\left(a^{\prime}\mid s^{\prime}\right)\right).$ (39)
@@
$\left\{\begin{array}[]{l}\sigma_{\text{mean }}^{c}<\vartheta_{c},\\ \log\left(\pi^{g}_{\phi}(a\mid s)<\kappa\right.,\\ n_{\text{steps }}>N_{g}.\end{array}\right.$ (40)

### v2

@@ II-D Overall Training Process
$J^{TD}_{\mathcal{Z}^{c}}(\zeta)=-\underset{\begin{subarray}{c}\left(s,a_{g},r,s^{\prime}\right)\sim\mathcal{B}_{g},a^{\prime}\sim\pi^{g}_{\bar{\phi}},\\ Z^{c}\left(s^{\prime},a^{\prime}\right)\sim\mathcal{Z}^{c}_{\bar{\zeta}}\left(\cdot\mid s^{\prime},a^{\prime}\right)\end{subarray}}{\mathbb{E}}\left[\log\mathcal{P}\left(y_{z}^{g}\mid\mathcal{Z}^{c}_{\zeta}(\cdot\mid s,a)\right)\right],$ (38)
@@
$y_{z}^{g}=r+\gamma\left(Z^{c}\left(s^{\prime},a^{\prime}\right)-\alpha\log\pi^{g}_{\bar{\phi}}\left(a^{\prime}\mid s^{\prime}\right)\right).$ (39)
@@
$\left\{\begin{array}[]{l}\sigma_{\text{mean }}^{c}<\vartheta_{c},\\ \log\left(\pi^{g}_{\phi}(a\mid s)\right)<\kappa,\\ n_{\text{steps }}>N_{g}.\end{array}\right.$ (40)